Two programmers started working on their computer programs at 9:00 am. The first programmer will take a Uniform(1hr,4hrs) time to finish the program.The second programmer will need a gamma(5hrs,1hr) time. Which programmer has a better chance to finish before 11:00am?


1
Expert's answer
2021-08-12T13:15:20-0400

We need to compute "P(X<2)" and "P(Y<2)" for a "Uniform(a=1\ hr, b=4\ hrs)" variable "X" and a "Gamma(r=5, \lambda=1\ hrs^{-1})" variable "Y" and compare them.

For the Uniform distribution with density "f(x)=\dfrac{1}{4-1}=\dfrac{1}{3}" for "1<x<4,"


"P(X<2)=\displaystyle\int_{-\infin}^{2}f(x)dx=\displaystyle\int_{1}^{2}\dfrac{1}{4-1}dx"

"=\big[\dfrac{1}{3}x\big]\begin{matrix}
 2 \\
 1
\end{matrix}=\dfrac{1}{3}\approx0.33333"

http://www.math.ntu.edu.tw/~hchen/teaching/StatInference/notes/lecture17.pdf

Gamma-Poisson relationship

There is a relationship between the gamma and Poisson distributions. If "Y" is a "Gamma(r, \lambda)" random variable, where "\alpha" is an integer, then for any "y,"


"P(Y\leq y)=P(Z\geq r),"

where "Z\sim Po(y/\lambda)."

For the second programmer, use the Gamma-Poisson formula with a "Gamma(r=5, \lambda=1)" variable "Y" and a Poisson "( y/\lambda=2/1=2)" variable "Z,"


"P(Y<2)=P(Z\geq r)=P(Z\geq5)\approx 0.05265"

from the Poisson distribution table with parameter 2.


Thus, the first programmer has a better chance to finish before 11:00 am.